An I.Q test was administered to 5 persons before and after they were trained. The
results are given below :
I II III IV V
I.Q before training 110 120 123 132 125
I.Q after training 120 118 125 136 121
Test whether there is any change in I.Q. after the training programme
(Given t0.01(4) = 4.6)

Grade:12

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
The results are given below : I II III IV V I.Q before training 110 120 123 132 125
I.Q after training 120 118 125 136 121
deviation in each case is 10 -2 2 4 -4
sum of d=10
d2=100+4+4+16+16=140
d=sum of d/n=10/2=5 and s=sqrt((d2-nd)/n-1)=sqrt(140-5(4))/5-1=5.673
Hence there is no change of IQ after training since givent0.01(4) = 4.6
M Mirdushi
2 years ago
The average marks scored by 32 biys is 72 with SD of 8, while that for 36 girls is 70 with SD of 6. Test at 1% level whether biys perform better than girls.